OpenStudy (anonymous):

Check @jim_thompson5910 In the right triangle below, a = 6 cm and b = 8 cm. What is c? A. 10 cm because c ^2 = a^ 2 + b ^2 <----- B. 14 cm because c = a + b C. 48 cm because c = ab D. 100 cm because c ^2 = a ^2 +b ^2

OpenStudy (fibonaccichick666):

Yup, but can You tell me what the name of the theorem you used was?

OpenStudy (anonymous):

that is correct

OpenStudy (baseballer2014):

\[a^2+b^2=c^2\] \[(6)^2+(8)^2=c^2\] \[36+64=c^2\]\[100=c^2 \] \[c=10\] it is the pythagorean theorem

OpenStudy (anonymous):

Yeah! I used the triangle thing theory. I can't pronounce it or spell it. :)

OpenStudy (baseballer2014):

It's the Pythagorean Theorem

OpenStudy (anonymous):

yes. :)

OpenStudy (fibonaccichick666):

alright, good, just wanted to make sure you knew